# Generating multipartite nonlocality to benchmark quantum computers

### v1

@@ I Introduction
To explain the proposed method, we will introduce graph states that can be readily prepared on a given connectivity. Accordingly, we discuss examples of Bell inequalities associated to graph states that show an exponential scaling of the fraction $D_{n}$ in Sec. II. Afterwards we will explain in Sec. III the method to measure the Bell inequalities. As we propose to evaluate the Bell inequalities by random sampling in Sec. III.2, we first formulate the Bell test as a hypothesis test in Sec. III.1. After we discuss the sample complexity in Sec. III.3, we will apply the method exemplarily to the Bell inequalities of the GHZ and LC state in Sec. IV. These Bell inequalities cover the extreme cases of connectivities in quantum computers. Finally, we use this to benchmark actual architectures in Sec. V and include a simulation for the IBM Eagle quantum processor in Sec. VI.
@@ II Bell inequalities with an exponential nonlocality
For the LC state in turn we consider the Bell inequalities in Ref. [30] that are defined in case the number of qubits is a multiple of three. The Bell operator is
@@
which takes the maximal value for the $n$-qubit LC state. As all terms in Eq. (7) are stabilizing operators of the LC state, we have $Q_{n}^{\text{LC}}=4^{n/3}$. The classical bound in turn is
@@ IV Analysis of the Bell inequalities for the GHZ and LC state
Figure 3: Necessary number of random observables $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ has a confidence of $\gamma=5\sigma$.
In this section, we are going to apply the method to the Bell inequalities for graph states. Especially, we will look at the Bell inequalities for the GHZ and the LC state. As described in the motivation, these Bell inequalities are promising to detect a large violation.
For this purpose, we express the observed expectation value as a fraction $\alpha$ of the quantum bound, i.e.,
^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}.$ (19)
As we have seen in Eq. (4), to observe a violation of the Bell inequality we have to have $\alpha>D_{n}^{-1}\overset{n\rightarrow\infty}{\rightarrow}0$. In case a violation is observed, it is $t=\alpha Q_{n}-C_{n}$. The number of random terms that are necessary to ensure a confidence level $\gamma$ is given by Eq. (15) that takes the form
We note that for the Bell inequalities in Sec. II, the number of terms equals the quantum bound, i.e., $M=Q_{n}$. As $D_{n}^{-1}$ converges to zero in the limit $n\rightarrow\infty$, we can conclude that $L$ converges against some constant value for fixed $\alpha$. This is shown for the Bell inequalities associated to the GHZ and LC state in Fig. 3.
For both the GHZ state as well as the LC state, we have assumed an observed violation of ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ with $\alpha=0.6$. This moreover implies that the assumed violation increases with the number of qubits $n$. We thus stress that Fig. 3 is only valid in case at least a value ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ is observed and does not show the generic scaling of $L$ with $n$. Exemplary we can make the observation:
Observation 1.
In case a violation ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ of the Bell inequality associated to the GHZ state or the LC state for $n=81$ qubits has been observed by sampling $L=80$ random terms, the result has a confidence level of $\gamma=5\sigma$.
@@
We note that for small $n$, $L$ exceeds the number of contexts of the Bell operator. This, however, is not a contradiction as we do not exclude that a term is sampled multiple times.
Finally, we see in Fig. 3 that already for relatively small $n$ the classical bound $C_{n}$ becomes negligible compared to the observed violation $0.6\times Q_{n}$. Therefore, it does not have an effect on $L$ for large $n$. We thus show the contours of the confidence levels $\gamma=1\sigma,2\sigma$ and $5\sigma$ in the limit $n\rightarrow\infty$ for a given observation $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ with $L$ random terms in Fig. 4.
V Bell nonlocality as benchmark
We now discuss Bell nonlocality as a benchmark for quantum computers. We will investigate the Bell violations that can be achieved on current quantum computers. For this purpose, we first consider the connectivity graphs of different quantum computers and discuss what Bell inequalities can be used. Afterwards, we include noise to assess realistic violations.
V.1 Connectivities of current quantum computers
As the optimal Bell inequality that is associated to the connectivity graph is in general hard to determine, we will focus on the Bell inequalities for the GHZ and the LC state. In the following, we are interested in the largest GHZ and LC state that can be natively prepared on the different layouts in Fig. 5, i.e., that can be prepared by the available two-qubit gates only.
@@ V.1 Connectivities of current quantum computers
As we only know the Bell inequality for the LC state with a number of qubits that is a multiple of three, we search for the longest path of length divisible by three. The longest paths that fulfill this restriction are drawn pink in Fig. 5. The connectivity graphs in Fig. 5a and Fig. 5b allow to prepare a $3$-qubit LC state. The largest simple path on the $20$-qubit full-connectivity graph in Fig. 5c with length divisible by three has length $18$. The quantum computer in Fig. 5c thus allows to check the Bell inequality for the $18$-qubit LC state. To find the longest path is an NP-complete problem [49]. We can thus not verify if the marked paths for the layouts in Fig. 5d and Fig. 5e are indeed the longest paths. In Fig. 5d, we have identified a $48$-qubit path as longest simple path. The layout in Fig. 5e in turn allows the preparation of the $108$-qubit LC state.
V.2 Noise
In the following section, we will discuss the effect of noise. The goal is to roughly estimate the violation that can be realistically observed on quantum computers. Commonly, the errors of quantum computers are specified in terms of the error rates for single-qubit gates, two-qubit gates and readout. The error rates for IBM’s Eagle processor ibm_brisbane and Google’s Sycamore processor are shown in Tab. 1. The error rate for the single-qubit gates is typically about an order smaller than the other error rates. We thus neglect the single-qubit errors. The readout error on the other side can be mitigated by classical post-processing [50, 51]. We thus focus on the error in the two-qubit gates. For this purpose, we will consider a simple depolarisation noise model.
@@ V.2 Noise
IBM Eagle $2.417\times 10^{-4}$ $7.409\times 10^{-3}$ $1.390\times 10^{-2}$
@@
Table 1: Error rates of IBM’s Eagle processor ibm_brisbane calibrated on 27.02.2024 18:32:14 [45] and Google’s Sycamore processor [13]. Google gives the error rates for the cases that the gates are performed isolated or simultaneously on all qubits.
@@
$\rho=p\outerproduct{G}{G}+(1-p)\frac{\mathbbm{1}}{2^{n}}.$ (21)
The probability that no error in the two-qubit gates occurs is $p=(1-p_{2})^{N_{2}}$, where $p_{2}$ is the error rate and $N_{2}$ the number of the two-qubit gates. The observed violation is thus
$\expectationvalue{\mathcal{B}}=p\times Q_{n}.$ (22)
@@
In Fig. 7, we show the circuits to prepare the LC and GHZ state in the case of $n=6$ qubits. For the LC state the CZ gates can be performed simultaneously and thus the circuit exhibits a constant depth of three independent of the number of qubits. For the GHZ state however the preparation with CZ gates requires the gates to be applied consecutively. The circuit depth thus grows with the number of qubits. The number of CZ gates, however, is equal.
@@
Table 2: Results for the depolarisation noise model that takes the 2-qubit gate error into account. For the noise data of Google’s Sycamore [13] and IBM’s Eagle processor [45] we show the violations in terms of the quantum bound $Q$. The last column shows the number of random terms $L$ that have to be sampled to verify the violation with a confidence of $5\sigma$.
@@
The results in Tab. 2 show that the simple noise model predicts a violation of more than $0.6\times Q$ for Google’s Sycamore processor and a bit less for ibm_brisbane. We note however that we considered more qubits on the IBM machine. A violation can thus be verified on both machines with $L\sim\mathcal{O}(10)-\mathcal{O}(10^{2})$. This is of the same order as in Sec. IV. However, we stress that the considered noise model does not take the 1-qubit gate errors and readout errors into account. Real quantum computers moreover do not implement all gates natively. IBM’s Eagle processor for example does not support the CZ gate. Rather, it has to be composed by the available gates. The circuit in practice thus contains more gates and possibly exhibits a larger depth. For these reasons the noise model probably overestimates the violation. However, it is still interesting to note that the noise model predicts the number of samples $L$ to increase exponentially in $n$. This is shown in Fig. 8.
Figure 8: Number of terms $L$ that have to be sampled to verify the violation predicted by the depolarisation noise model with a confidence level of $\gamma=5\sigma$.
VI Simulation for an IBM quantum computer
@@
Figure 10: Simulation for the IBM Eagle quantum processor. The simulation uses the error rates of the real device ibm_brisbane [45] and the violation is estimated by measuring $L=800$ random terms of the Bell inequality $K=1$ times each. In (a) the average expectation value of the Bell inequality $\mathcal{B}^{\text{LC}}$ is shown for the LC state, whereas (b) shows the violation of $\mathcal{B}^{\text{GHZ}}$ for the GHZ state. In both cases, the expectation values are averaged over $10$ repetitions and the error bars show the standard deviation.
@@
After the preparation, $L$ random terms of the corresponding Bell inequality are measured. The measurement of each random term is not repeated, i.e., $K=1$. Fig. 10 shows the average expectation values of the Bell inequalities for the LC and GHZ state of up to $n=24$ qubits. We have chosen $L=800$ random terms and the average is taken over $10$ repetitions. Moreover, the number of qubits is a multiple of three as only in this case a good Bell inequality for the LC state is known. Fig. 10 shows that for both states the simulation predicts a Bell violation that increases exponentially with $n$. The LC state however shows a slightly higher relative violation compared to the GHZ state. This can be seen in Fig. 11a. The plot in Fig. 11a displays the observed violation as a fraction of the total number of terms in the Bell inequality, i.e., $t^{2}/M^{2}=(\langle{\hat{\mathcal{B}}}\rangle-C_{n})^{2}/M^{2}$. We fit the data points to an exponential and obtain
$\begin{split}\left(\frac{t^{2}}{M^{2}}\right)_{\text{LC}}(n)=\exp(-0.10n-0.65),\\ \left(\frac{t^{2}}{M^{2}}\right)_{\text{GHZ}}(n)=\exp(-0.17n-0.54).\end{split}$ (23)
This affirms that the relative violation of the GHZ state decreases faster with $n$ compared to the LC state. We attribute this to the larger circuit depth that is required for the GHZ state. The preparation of the GHZ state is thus more affected by noise. The smaller relative violation is also the reason for the larger $p$-values for the GHZ state, which we present in Fig. 9. Except for the cases of $n=3,15$ the $p$-values of the LC state are smaller, which implies a higher significance of the observed violation.
Finally, we use Eq. (23) to extrapolate the violation to larger $n$. Hoeffding’s inequality in Eq. (17) yields for $K=1$ and a target value for the confidence $\gamma$:
@@
The number of necessary sampled terms $L$ is shown in Fig. 11b for the LC state and in Fig. 11c for the GHZ state. We first point out the different scaling of $L$ compared to Fig. 3. Whereas in Fig. 3 $L$ decreases with $n$ until a plateau is reached, Fig. 11 suggests that $L$ increases exponentially with $n$. This however can be attributed to fact that Fig. 3 has been obtained with the assumption that the relative violation stays constant with the number of qubits $n$. But Fig. 11a indicates that this assumption is not true for a real quantum computer. However, we still obtain that the number of terms $L$ that have to be sampled is much smaller than the total number of terms $M$.
As an example, we discuss $L$ for the case of $n=81$ qubits and a target confidence of $\gamma=5\sigma$:
$\begin{split}L_{\text{LC}}(n=81,\gamma=5\sigma)&=182813,\\ L_{\text{GHZ}}(n=81,\gamma=5\sigma)&=47170845.\end{split}$ (25)
The above numbers are much smaller than $M_{\text{LC}}\approx 10^{16}$ and $M_{\text{GHZ}}\approx 10^{24}$, but larger than the values predicted in Sec. IV. As previous, we attribute this to the unrealistic assumption that independent of $n$ a violation $\langle{\mathcal{B}_{n}}\rangle=0.6\times Q_{n}$ can be observed. For $n=24$ qubits the simulation yields ^$\langle{\hat{\mathcal{B}}^{\text{LC}}_{n}}\rangle/Q^{\text{LC}}_{n}=0.28$ for the LC state and ^$\langle{\hat{\mathcal{B}}^{\text{GHZ}}_{n}}\rangle/Q^{\text{GHZ}}_{n}=0.09$ for the GHZ state, which is already much smaller than $0.6$ in both cases.
@@
Figure 11: (a) Observed violation in relation to the number of terms in the Bell inequality, i.e., $t^{2}/M^{2}=(\langle{\hat{\mathcal{B}}}\rangle-C_{n})^{2}/M^{2}$. The data points are fitted by an exponential. From the fitted function $t^{2}/M^{2}$ we estimate the number of random terms $L$ that have to be sampled to reach a confidence of $\gamma=1\sigma,2\sigma,5\sigma$. (b) shows the results for the LC state, whereas (c) illustrates the predictions for the GHZ state.
VII Conclusion

### v2

@@ I Introduction
To explain the proposed method, we will introduce graph states that can be readily prepared on a given connectivity. Accordingly, we discuss examples of Bell inequalities associated to graph states that show an exponential scaling of the fraction $D_{n}$ in Sec. II. Afterwards we will explain in Sec. III the method to measure the Bell inequalities. As we propose to evaluate the Bell inequalities by random sampling in Sec. III.2, we first formulate the Bell test as a hypothesis test in Sec. III.1. After we discuss the sample complexity in Sec. III.3, we will apply the method exemplarily to the Bell inequalities of the GHZ and LC state in Sec. IV. These Bell inequalities cover the extreme cases of connectivities in quantum computers. In doing so, we consider actual architectures of current quantum computers. Finally, we include a simulation for the IBM Eagle quantum processor in Sec. V.
@@ II Bell inequalities with an exponential nonlocality
For the LC state, in turn, we consider the Bell inequalities in Ref. [30] that are defined in case the number of qubits is a multiple of three. The Bell operator is
@@
which takes the maximal value for the $n$-qubit LC state. As all terms in Eq. (7) are stabilizing operators of the LC state, we have $Q_{n}^{\text{LC}}=4^{n/3}$. The classical bound, in turn, is
@@ III.3 Number of measurement repetitions
Figure 3: Necessary number of random observables $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ has a confidence of $\gamma=5\sigma$. $L$ is plotted as a function of the number of qubits $n$.
As an example, we consider the case that a violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ is observed. This value is motivated by Ref. [36], where they prepared a $51$ qubit LC state with fidelity of approximately $0.6$. The necessary number of mesurement settings $L$ is shown in Fig. 3. We note that for small $n$, $L$ exceeds the number of contexts of the Bell operator. This, however, is not a contradiction as we do not exclude that a term is sampled multiple times. Exemplary we can make the observation:
Observation 1.
In case a violation ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=0.6\times Q_{n}$ of the Bell inequality associated to the GHZ state or the LC state for $n=51$ qubits has been observed by sampling $L=80$ random terms, the result has a confidence level of $\gamma=5\sigma$.
@@ IV Analysis of the Bell inequalities for the GHZ and LC state
In this section, we are going to apply the method to the Bell inequalities for graph states. Especially, we will look at the Bell inequalities for the GHZ and the LC state. As described in the motivation, these Bell inequalities are promising to detect a large violation. The discussion in the previous section has shown that this is necessary to verify the violation from few measurement settings with high significance. We start with the question how the LC and GHZ state can be prepared on a quantum computer with a given two-qubit connectivity. Afterwards, we consider the effect of noise on the preparation with a simple depolarisation noise model, which gives insight into the sample complexity of the method.
IV.1 Connectivities of current quantum computers
@@
As the optimal Bell inequality that is associated to the connectivity graph is in general hard to determine, we will focus on the Bell inequalities for the GHZ and the LC state. In the following, we are interested in the largest GHZ and LC state that can be prepared on the different layouts in Fig. 5. Ideally, the preparation only requires the set of basis gates that are directly implemented by the quantum computer.
@@
As we only know the Bell inequality for the LC state with a number of qubits that is a multiple of three, we search for the longest path of length divisible by three. The longest paths that fulfill this restriction are drawn pink in Fig. 5. The connectivity graphs in Fig. 5a and Fig. 5b allow to prepare a $3$-qubit LC state. The largest simple path on the $20$-qubit full-connectivity graph in Fig. 5c with length divisible by three has length $18$. The quantum computer in Fig. 5c thus allows to check the Bell inequality for the $18$-qubit LC state. To find the longest path is an NP-complete problem [49]. We can thus not verify if the marked paths for the layouts in Fig. 5d and Fig. 5e are indeed the longest paths. In Fig. 5d, we have identified a $48$-qubit path as longest simple path. The layout in Fig. 5e, in turn, allows the preparation of the $108$-qubit LC state.
The use of the LC state corresponding to the longest single path implies that we are not using all the qubits of the quantum computer. Of course, if one focuses on a single graph state that only covers part of the qubits, then one obtains only partial information about the computer. However, in principle, one can cover all the qubits by LC and GHZ states and, in this way, obtain more information. The quality of the benchmark depends on the effort and detail one wants to invest in it. But, even with a moderate effort, our method allows making simple statements such as quantum computer $A$ manages to produce a value $\langle{\hat{\mathcal{B}}}\rangle$ of the Bell parameter with $N$ qubits in a LC state, which are valuable to compare $A$ to other computers.
IV.2 Noise
In the following section, we will discuss the effect of noise. The goal is to roughly estimate the violation that can be realistically observed on quantum computers and how the violation scales with the number of qubits $n$. Commonly, the errors of quantum computers are specified in terms of the error rates for single-qubit gates, two-qubit gates and readout. The average error rates for IBM’s Eagle processor ibm_brisbane and Google’s Sycamore processor are shown in Table 1. The error rate for the single-qubit gates is typically about an order smaller than the other error rates. The readout error on the other side can be mitigated by classical post-processing [50, 51]. To consider the effect of noise on the Bell violation, we will consider a simple depolarisation noise model.
@@ IV.2 Noise
IBM Eagle $4.322\times 10^{-4}$ $1.019\times 10^{-2}$ $2.434\times 10^{-2}$
@@
Table 1: Average error rates of IBM’s Eagle processor ibm_brisbane calibrated on 27.02.2024 18:32:14 [45] and Google’s Sycamore processor [13]. Google gives the error rates for the cases that the gates are performed isolated or simultaneously on all qubits. The error rates are averaged over all gates or the readout error of all qubits.
@@
$\rho=\alpha\outerproduct{G}{G}+(1-\alpha)\frac{\mathbbm{1}}{2^{n}}.$ (19)
The probability that no error in the preparation occurs is $\alpha=(1-p_{1})^{N_{1}}(1-p_{2})^{N_{2}}(1-p_{r})^{n}$, where $p_{1}$ and $p_{2}$ are the average error rates for single- and two-qubit gates and $p_{r}$ is the average readout error. $N_{1}$ denotes the number of single-qubit gates. We note that idling qubits can be seen as an identity gate acting on the qubits. Identity gates are single-qubit gates and the error can also be described by the average error in the single-qubit gates [45]. $N_{2}$, in turn, is the number of two-qubit gates and $n$ the number of qubits that are prepared. The observed violation is thus
$\expectationvalue{\mathcal{B}}=\alpha\times Q_{n}.$ (20)
@@
In Fig. 6, we show the circuits to prepare the LC and GHZ state in the case of $n=6$ qubits. For the LC state the CZ gates can be performed simultaneously and thus the circuit exhibits a constant depth of three independent of the number of qubits. The number of single-qubit gates is $N_{1}=n+2$, which includes two identity operations for the idling qubits. There are in total $n-1$ CZ gates, such that $N_{2}=n-1$. For the GHZ state, however, the preparation with CZ gates requires the gates to be applied consecutively. The circuit depth thus grows with the number of qubits. In each step, $n-2$ of the qubits are idling. Therefore, there are in total $(n-1)(n-2)$ identity operations, such that the preparation requires in total $N_{1}=n+(n-1)(n-2)$ single qubit gates. The number of CZ gates is equal, i.e., $N_{2}=n-1$.
@@
Table 2: Results for the depolarisation noise model. For the noise data of Google’s Sycamore [13] and IBM’s Eagle processor [45] we show the violations in terms of the quantum bound $Q$. The last column shows the number of random terms $L$ that have to be sampled to verify the violation with a confidence of $5\sigma$.
@@
The results in Table 2 show that the simple noise model predicts a violation of approximately $0.1\times Q$ for Google’s Sycamore processor and around $0.02\times Q$ for ibm_brisbane. We note however that we considered more qubits on the IBM machine. Moreover, the violation of the GHZ state is smaller compared to the LC state. This indicates that the GHZ state is more affected by noise. A violation can accordingly be verified with $L\sim 2000$ measurement settings on Google’s Sycamore processor. On the IBM Eagle processor, in turn, $L\sim 40000$ settings are needed to verify nonlocality in the LC state and $L\sim 100000$ for the GHZ state. Real quantum computers, moreover, do not implement all gates natively. IBM’s Eagle processor for example does not support the CZ gate. Rather, it has to be composed by the available gates. The circuit in practice thus contains more gates and possibly exhibits a larger depth. For these reasons the noise model overestimates the violation. However, the noise model is still useful to assess the scaling of the sample complexity.
IV.3 Sample complexity
Figure 7: Necessary number of random observables $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ with $\alpha_{\text{GHZ}}=\exp(-an^{2})$ or $\alpha_{\text{LC}}=\exp(-bn)$ has a confidence of $\gamma=5\sigma$. $L$ is plotted as a function of the number of qubits $n$.
In the previous section, we have seen that noise causes the Bell violation to decline exponentially with the number of qubits, i.e., in leading order we have
$\alpha_{\text{GHZ}}$ $=\exp(-an^{2}),$ (21a)
$\alpha_{\text{LC}}$ $=\exp(-bn)$ (21b)
with $a,b>0$. As we have seen in Eq. (4), a violation of the Bell inequality can still be observed for large $n$ if $\alpha$ vanishes slower than the fraction $D_{n}^{-1}$, i.e., $\alpha>D_{n}^{-1}\overset{n\rightarrow\infty}{\rightarrow}0$. We note that for the LC state this is the case for $b<\log(2)/3$. In this parameter regime, the relative violation ^$\langle{\hat{\mathcal{B}}}\rangle/C$ of the Bell inequality for the LC state is still increasing with $n$. For the GHZ state, however, the scaling in Eq. (21) shows that the effect of noise increases faster than the quantum bound. A violation of the Bell inequality for the GHZ state can thus only be observed for $n<\ln(2)/(4a)+\sqrt{(\ln(2)/(4a))^{2}-\ln(2)/a}$. In case a violation is observed, it is $t=\alpha Q_{n}-C_{n}$. The number of random terms that is necessary to ensure a confidence level $\gamma$ is given by Eq. (15) and takes the form
We note that, for the Bell inequalities in Sec. II, the number of terms equals the quantum bound, i.e., $M=Q_{n}$. In case $\alpha$ also decreases exponentially, the number of random settings $L$ increases exponentially. This is shown in Fig. 7.
For both the GHZ state as well as the LC state, we have plotted $L$ as a function of the number of qubits $n$. The scaling for the GHZ state is due to the number of single-qubit gates. We therefore choose the parameter $a$ of the same order as the error in the single-qubit gates, i.e., $a=0.001,0.0005$ and $0.0001$. For the LC state in contrast all errors contribute to the leading term and we choose $b=0.1,0.01$ and $0.001$. Fig. 7 shows that for $a\leq 0.0001$ or $b\leq 0.01$ the number of required measurements are feasible on current quantum computers.
Finally, if the classical bound $C_{n}$ is negligible compared to the observed violation, it does not have an effect on $L$. We thus show the contours of the confidence levels $\gamma=1\sigma,2\sigma$ and $5\sigma$ in this limit for a given observation $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ with $L$ random terms in Fig. 8.
V Simulation for an IBM quantum computer
Figure 9: Simulation for the IBM Eagle quantum processor. The simulation uses the error rates of the real device ibm_brisbane [45] and the violation is estimated by measuring $L=800$ random terms of the Bell inequality $K=1$ times each. In (a), the average expectation value of the Bell inequality $\mathcal{B}^{\text{LC}}$ is shown for the LC state, whereas (b) shows the violation of $\mathcal{B}^{\text{GHZ}}$ for the GHZ state. In both cases, the expectation values are averaged over $10$ repetitions and the error bars show the standard deviation.
@@ V Simulation for an IBM quantum computer
After the preparation, $L$ random terms of the corresponding Bell inequality are measured. The measurement of each random term is not repeated, i.e., $K=1$. Fig. 9 shows the average expectation values of the Bell inequalities for the LC and GHZ state of up to $n=24$ qubits. We have chosen $L=800$ random terms and the average is taken over $10$ repetitions. Moreover, the number of qubits is a multiple of three as only in this case a good Bell inequality for the LC state is known. Fig. 9 shows that for both states the simulation predicts a Bell violation that increases exponentially with $n$. The LC state however shows a slightly higher relative violation compared to the GHZ state. This can be seen in Fig. 11a. The plot in Fig. 11a displays the observed expectation value as a fraction of the quantum bound, i.e., ^$\langle{\hat{\mathcal{B}}}\rangle/Q$. In agreement with the scaling that is predicted by the depolarisation noise model, we fit the logarithmic data to a linear function for the LC state and a quadratic function for the GHZ state. The fits yield
^$\frac{\langle{\hat{\mathcal{B}}_{n}^{\text{LC}}}\rangle}{Q_{n}^{\text{LC}}}$ $=\exp(-0.078n+0.248),$ (23a)
^$\frac{\langle{\hat{\mathcal{B}}_{n}^{\text{GHZ}}}\rangle}{Q_{n}^{\text{GHZ}}}$ $=\exp(-0.001n^{2}-0.078n+0.154).$ (23b)
This affirms that the relative violation of the GHZ state decreases faster with $n$ compared to the LC state. We attribute this to the larger circuit depth that is required for the GHZ state. The preparation of the GHZ state is thus more affected by noise. The smaller relative violation is also the reason for the larger $p$-values for the GHZ state, which we present in Fig. 10. Except for the cases of $n=3,15$ the $p$-values of the LC state are smaller, which implies a higher significance of the observed violation. Overall, the $p$-values show a large variation. We note that the $p$-values depend on the observed violation. Apparently, for some numbers of qubits $n$ the evaluation of the Bell inequalities is more affected by noise. We attribute this to the mapping to the actual architecture. For different sizes of the state, the optimal choice of qubits might be differently affected by noise. This also explains the variation of the error bars in Fig.  9. In case the evaluation is more affected by noise, not only is the observed violation smaller but it also appears reasonable that the value fluctuates more. In addition, we note that the error bars in Fig. 11a do not match the fitted exponential. This is because the plotted error bars show the standard deviation of the repeated simulations and do not cover all uncertainties. For example, the error bars do not include the uncertainties due to the different mappings of the quantum circuits to the actual architecture.
Finally, we use Eq. (23) to extrapolate the violation to larger $n$. We note that the circuits have to be adapted to the architecture of the quantum computer. For this task, we use the automatic transpilation provided in Qiskit, which involves an optimization such that the qubits are chosen that are least affected of noise. The extrapolation overestimates the violation for large $n$, since in this case such a choice is not possible anymore. Hoeffding’s inequality in Eq. (17) yields for $K=1$ and a target value for the confidence $\gamma$:
@@
The number of necessary sampled terms $L$ is shown in Fig. 11b for the LC state and in Fig. 11c for the GHZ state.
@@
Figure 11: (a) Observed violation as a fraction of the quantum bound, i.e., ^$\langle{\hat{\mathcal{B}}}\rangle/Q$. The logarithmic data is fitted by a linear function for the LC state and a quadratic function for the GHZ state. From the fitted function, we estimate the number of random terms $L$ that have to be sampled to reach a confidence of $\gamma=1\sigma,2\sigma,5\sigma$. (b) shows the results for the LC state, whereas (c) illustrates the predictions for the GHZ state.
As an example, we discuss $L$ for the case of $n=108$ qubit LC state, i.e., the largest simple path for the architecture. For better comparison, we also calculate $L$ for the $n=108$ qubit GHZ state. To reach a target confidence of $\gamma=5\sigma$ the values are:
$L_{\text{LC}}(n=108,\gamma=5\sigma)$ $=330997173,$ (25a)
$L_{\text{GHZ}}(n=108,\gamma=5\sigma)$ $=1.825\times 10^{23}.$ (25b)
Nonlocality of the LC state can be verified by a large number of measurement settings. For the GHZ state, in turn, the extrapolation predicts also a violation. To verify this violation, however, requires an infeasible number of measurements. This shows again that the GHZ state is more affected by noise than the LC state and that the LC state is thus more promising to detect large multipartite nonlocality. The above numbers are still smaller than $M_{\text{LC}}\approx 10^{21}$ and $M_{\text{GHZ}}\approx 10^{32}$, but much larger than the values predicted in Sec. IV. We attribute this to the fact that the generic depolarisation noise model does not cover all the noise in a specific quantum computer. Recall, however, that our main goal is to detect large-scale quantum nonlocality and our method can still be used to assess the confidence of an observed Bell violation. Of course, the verification with a certain confidence may require more measurements for specific computers. In the worst case (e.g., when $\alpha$ drops and the $p$-value is not sufficient), one just needs to make additional measurements.
VI Conclusion